OpenStudy (help!!!!):

16. In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; mathematics is worth 3 points; History is worth 2 points; and English is worth 3 points. The values of the grade letters are as follows, A= 4, B=3, C=2, D=1, F=0. What is the GPA of a student who made a “C” in Trigonometry, a “B” in American History, an “A” in Botany, and a “B” in Microbiology? A. 2.59 B. 2.86 C. 3.08 D. 3.33 E. 3.67

OpenStudy (help!!!!):

im clueless

OpenStudy (help!!!!):

so 2/3+3/2+4/4+4/4?

OpenStudy (kinggeorge):

This question is basically a question on averages. You add up the values*points, and then divide by points. For example, suppose there was only a “C” in Trigonometry, a “B” in Microbiology. Trigonometry was worth 3 points, and a C has a value of 2. M.Biology was worth 4 points, and a B has a value of 3. So the GPA would be\[\frac{3*2+4*3}{3+4}=\frac{18}{7}\]Does this help?

OpenStudy (help!!!!):

yeah

OpenStudy (kinggeorge):

Can you show me what the setup would be for the whole problem?

OpenStudy (help!!!!):

so 2*3+3*2+4*4+4*3/3+2+4+4

OpenStudy (kinggeorge):

Looks great.

OpenStudy (help!!!!):

40/13 =3.08

OpenStudy (kinggeorge):

perfect.

OpenStudy (help!!!!):

is there an equation for this?